# Deep Unrolled Meta-Learning for Multi-Coil and Multi-Modality MRI with Adaptive Optimization

### v1

@@ Abstract
We propose a unified deep meta-learning framework for accelerated magnetic resonance imaging (MRI) that jointly addresses multi-coil reconstruction and cross-modality synthesis. Motivated by the limitations of conventional methods in handling undersampled data and missing modalities, our approach unrolls a provably convergent optimization algorithm into a structured neural network architecture. Each phase of the network mimics a step of an adaptive forward-backward scheme with extrapolation, enabling the model to incorporate both data fidelity and nonconvex regularization in a principled manner. To enhance generalization across different acquisition settings, we integrate meta-learning, which enables the model to rapidly adapt to unseen sampling patterns and modality combinations using task-specific meta-knowledge. The proposed method is evaluated on the open source datasets, showing significant improvements in PSNR and SSIM over conventional supervised learning, especially under aggressive undersampling and domain shifts. Our results demonstrate the synergy of unrolled optimization, task-aware meta-learning, and modality fusion, offering a scalable and generalizable solution for real-world clinical MRI reconstruction.
@@
1 Motivation
Magnetic Resonance Imaging (MRI) is one of the most advanced and widely used medical imaging techniques due to its non-invasive nature and superior soft-tissue contrast. Among various MRI techniques, Parallel MRI (pMRI) has emerged as a state-of-the-art method in clinical settings. pMRI systems employ an array of multiple receiver coils that simultaneously acquire k-space (Fourier domain) measurements, thereby enabling faster data acquisition and enhanced spatial resolution. This parallel acquisition strategy makes pMRI especially valuable in time-sensitive or high-resolution imaging scenarios.
Mathematically, consider a pMRI system with $N_{c}$ receiver coils acquiring 2D MR images at resolution $m\times n$ (we treat a 2D image $\mathbf{v}\in\mathbb{C}^{m\times n}$ and its column vector form $\mathbf{v}\in\mathbb{C}^{mn}$ interchangeably hereafter). Let $\mathbf{P}\in\mathbb{R}^{p\times mn}$ be the binary matrix representing the undersampling mask with $p$ sampled locations in k-space, $\mathbf{s}_{i}\in\mathbb{C}^{mn}$ the coil sensitivity, and $\mathbf{f}_{i}\in\mathbb{C}^{p}$ the partial k-space data at the $i$th receiver coil for $i=1,\dots,N_{c}$. Therefore $\mathbf{f}_{i}$ and the image $\mathbf{v}$ are related by $\mathbf{f}_{i}=\mathbf{P}\mathbf{F}(\mathbf{s}_{i}\cdot\mathbf{v})+\mathbf{n}_{i}$, $\mathbf{F}\in\mathbb{C}^{mn\times mn}$ stands for the discrete Fourier transform, and $\mathbf{n}_{i}$ is the unknown acquisition noise. Then an image space reconstruction methods can generally be formulated as an optimization problem as follows:
$\min_{\mathbf{v}}\ \sum\nolimits^{N_{c}}_{i=1}\frac{1}{2}\|\mathbf{P}\mathbf{F}(\mathbf{s}_{i}\cdot\mathbf{v})-\mathbf{f}_{i}\|^{2}+R(\mathbf{v}),$ (1)
However, two central challenges arise when solving this problem [1]. First, the design and tuning of the regularizer $R(\mathbf{v})$ (including the regularization weight).including its functional form and regularization weight—are critical to achieving high-quality reconstruction. Inappropriate choices may lead to over-smoothing or artifact amplification. Second, accurate knowledge of the coil sensitivity maps $\{\mathbf{s}_{i}\}$. Although some recent techniques attempt to circumvent explicit sensitivity estimation (e.g., self-calibrating methods), the absence of precise coil sensitivities can still degrade reconstruction quality, especially in highly accelerated settings.
This paper presents a deep learning-based approach for Magnetic Resonance Imaging (MRI) reconstruction, as proposed in [2]. The method addresses key challenges in parallel MRI (pMRI) by jointly reconstructing all channel-wise coil images through a unified regularization strategy. Instead of applying regularization independently to each coil image, the framework first learns a combination operator that fuses multi-coil data into a single, composite image. Regularization is then applied to this combined image, leveraging the rich spatial and frequency information available across all channels. The full field-of-view (FOV) image is thus synthesized iteratively through the learned combination operator [3], which is refined at each network phase.
To further enhance reconstruction quality, the method alternates between updating coil images in the spatial (image) domain and refining their representations in the frequency (k-space) domain. This bidirectional update mechanism ensures both data consistency and structural integrity.
Given that pMRI data are inherently complex-valued and phase information is diagnostically significant, the network incorporates complex-valued convolutions and nonlinear activations to model both magnitude and phase accurately. Additionally, the initial step of the reconstruction pipeline interpolates the undersampled k-space data to generate a high-quality starting point for the iterative procedure.
The overall approach unrolls an optimization algorithm into a structured convolutional neural network (CNN), where each network phase corresponds to one iteration of the underlying algorithm. Experimental evaluations demonstrate that this method achieves superior performance compared to several state-of-the-art pMRI reconstruction techniques, validating the efficacy of the proposed architecture.
@@ 2 Introduction
MRI is one of the most prominent non-invasive and non-ionizing medical imaging technologies to generate precise in-vivo issue images for disease diagnosis and medical analysis. MR examination can provide images of different contrasts (modalities), such as inversion recovery attenuated with T1 , inversion attenuated with T2 , and inversion recovery attenuated with T2 fluid (Flair). Capturing images of an anatomy with multi-modality acquisitions enhance the diagnostic information since each modality captures specific physical characteristic of the tissues under the same anatomy. Radiologists can locate subtle lesions and the extent of their deterioration, such as a developing tumor, by comparing multiple contrast MRI[4]. However, collecting multiple MR modalities is often challenging due to the long scanning time, patient movement, and corruption caused by artifacts. This paper developed new techniques that integrate the variation method with the deep learning approach for multi-modality magnetic resonance image (MRI) synthesis and reconstruction simultaneously.
Medical image synthesis has been proposed as an effective solution, where missing modalities are synthesized from existing ones. However, most state-of-the-art image synthesis methods are converting true available modalities to another. The major drawback is that the input images are acquired from a fully scanned k-space measurement, which requires excessive acquisition times to acquire true MR data.
In this project, instead of acquiring the fully scanned MR data in the k-space, they focus on missing modality synthesis by inputting the partial (undersampled) k-space measurement to accelerate the acquisition process. In addition, in this project they propose a convergence guaranteed algorithm with rigorous proves. They consider to conduct a learning based convolutional neural network (CNN) to simultaneously learn the fusion operator from each available modality to obtain the missing modality and reconstruct all modality images jointly by the following model:
@@ 3 Related Work
More recently, a diffusion model for accelerated MRI[20] and quantitative MRI [21]reconstruction using domain-conditioned prior guidance. This method leverages generative modeling to learn sample-consistent priors conditioned on domain information, achieving state-of-the-art results in both structural fidelity and quantitative accuracy. Bian et al. [22] introduced a diffusion-based generative framework for accelerated MRI and quantitative MRI (qMRI) reconstruction. Their approach leverages domain-conditioned prior guidance, where learned priors are tailored to specific anatomical regions or acquisition protocols.
@@ 4 Method
To solve the multi-modality and multi-coil MRI reconstruction problem under undersampled measurements, we propose an adaptive forward-backward optimization algorithm with extrapolation, which is integrated into the deep network structure through algorithm unrolling. The objective is to minimize a composite function $\phi_{\eta}(x)=f(x)+r_{\eta}(x)$, where $f$ is the data fidelity term and $r_{\eta}$ is a smoothed nonconvex regularizer with parameter $\eta$.
The algorithm follows a two-stage update with an adaptive learning rate for $\eta$, inspired by proximal gradient methods and extrapolated momentum schemes. The detailed steps are described in Algorithm 1.
1:  Input: Initial point $x_{0}$, parameters $10<\sigma,\tau_{0},\mu,\delta<$ 1, $\eta_{0}>0$, tolerance $\epsilon_{\text{tol}}>0$
2:  for $k=0,1,\dots,K$ do
3:     Perform a forward step using the gradient of the data fidelity term:
$z_{k}=x_{k}-\alpha_{k}\nabla f(x_{k})$
4:     Perform a backward (proximal-like) step on the regularization term:
$x_{k+1}=\begin{cases}u_{k}=z_{k}-\beta_{k}\nabla r_{\eta_{k}}(z_{k}),&\text{if}\phi_{\eta_{k}}(u_{k})\leq\phi_{\eta_{k}}(x_{k})-\frac{\delta}{2}\|u_{k}-x_{k}\|^{2}\\ v_{k}=x_{k}-\gamma_{k}\nabla\phi_{\eta_{k}}(x_{k}),&\text{otherwise}\end{cases}$
This condition ensures sufficient decrease of the objective value or defaults to a full gradient descent step.
5:     Apply extrapolation to accelerate convergence:
$y_{k}=x_{k+1}+\tau_{k}(x_{k+1}-x_{k})$
6:     Accept the extrapolated point if it further reduces the objective:
$x_{k+2}=\begin{cases}y_{k},&\text{if }\phi_{\eta_{k}}(y_{k})\leq\phi_{\eta_{k}}(x_{k+1})\text{ and set }\tau_{k+1}=\min\left\{\frac{\tau_{k}}{\mu},1\right\}\\ x_{k+1},&\text{otherwise, set }\tau_{k+1}=\mu\tau_{k}\end{cases}$
This adaptive control of $\tau_{k}$ ensures stability in case of divergence.
7:     Update the smoothing parameter $\eta_{k}$ if sufficient gradient decay is observed:
$\text{If }\|\nabla\phi_{\eta_{k}}(x_{k+1})\|<\sigma\eta_{k},\text{ then }\eta_{k+1}=\sigma\eta_{k};\text{ else }\eta_{k+1}=\eta_{k}$
8:     if $\sigma\eta_{k}<\epsilon_{\text{tol}}$ then
9:        Terminate
10:     end if
11:  end for
12:  Output: $x_{K}$
Algorithm 1 Adaptive Forward-Backward Algorithm with Extrapolation
@@
Loss function is defined as
$\ell(x_{0},x_{1};\hat{x}_{0},\hat{x}_{1})=\|x_{0}-\hat{x}_{0}\|^{2}+\lambda_{1}(1-ssim(x_{0},\hat{x}_{0}))+\lambda_{2}\|\tilde{\mathbf{g}}_{0}(\mathcal{G}(\mathbf{g}_{1}(x_{1})))-\hat{x}_{0}\|^{2}+\lambda_{3}\|x_{1}-\hat{x}_{1}\|^{2},$ (2a)
In this formulation, $x_{0}$ denotes synthesised image and $x_{1}$ corresponds to the reconstructed image produced at the final (i.e., $T$-th) phase of the network. The terms $\hat{x}_{0}$ and $\hat{x}_{1}$ represents their respective ground-truth of $x_{0}$ and $x_{1}$. represent their respective ground-truth references. The loss integrates four components: (i) a pixel-wise reconstruction error for the synthesized image, (ii) a structural similarity (SSIM) term to promote perceptual fidelity[23], (iii) a consistency loss through a composed mapping of the reconstructed image back to the synthesized domain, and (iv) a direct reconstruction loss between the final output and its ground-truth.
5 Experiment
The test dataset used in our experiments is sourced from the publicly available fastMRI dataset [24], which provides 15-channel undersampled k-space data. For our evaluation, we focus on Cartesian equispaced undersampling masks with acceleration factors of 4-fold and 8-fold. After careful verification, we confirmed that the 4-fold acceleration mask used in fastMRI is identical to the sampling pattern adopted in our training and testing phases. Due to the large volume of available test data, we randomly selected 10 different sequences, and from each sequence, we extracted 3 successive slices of multi-coil k-space data. This results in a total of 30 multi-coil MR slices used for evaluation.
The reconstruction model was tested using weights obtained from the fifth unrolled phase of the network (Phase 5). The quantitative results achieved on the selected 30 test samples demonstrate the effectiveness of our method: the average Peak Signal-to-Noise Ratio (PSNR) is 41.7114 dB, the average relative error is 0.026917, and the average Structural Similarity Index Measure (SSIM) is 0.9719. These values reflect a high-fidelity reconstruction with minimal information loss.
It is important to note that the evaluation metrics were computed using our in-house quantifier implementation as described in the methodology section, and not with the official fastMRI challenge leaderboard metrics, due to limitations in submitting custom test sets to the official evaluation platform. Nevertheless, our metric computation follows standard definitions of PSNR, relative error, and SSIM, ensuring fair and reproducible comparison.
This evaluation strategy was chosen to strike a balance between rigorous testing and computational feasibility. Although we only tested a subset of the full fastMRI dataset, the randomly selected sequences and slices offer representative insights into the performance of the model under typical clinical variations in anatomy and scan geometry. In future work, we aim to scale this testing pipeline to encompass a larger and more diverse portion of the dataset, and to benchmark performance under more challenging acceleration scenarios (e.g., 8x Cartesian or non-Cartesian undersampling).
To further validate the proposed method, we conducted experiments using data from the Multimodal Brain Tumor Segmentation Challenge 2018 (BraTS) [25]. We focused specifically on high-grade glioma (HGG) cases and used two imaging modalities: T1-weighted and T2-weighted MRIs. From each modality, we selected 30 patients and extracted 10 consecutive slices from the center of each volume, resulting in 300 images for training. For validation, we randomly selected 15 different patients, and an additional 6 patients (excluded from both training and validation) were used for testing, giving us 150 validation images and 60 test images.
We compared our meta-learning approach with conventional supervised learning. In the conventional setting, the network was trained on a single task at a time and optimized only the main network parameters $\theta$. In contrast, the meta-learning model trained across tasks and additionally learned task-specific meta-knowledge $\omega_{i}$. Both models used the same unrolled structure with 9 iterative phases (Algorithm 1). For conventional supervised learning, we merged the training and validation datasets into a unified set of 450 images to match the data availability.
Experimental results demonstrate that our meta-learning method consistently outperforms the conventional approach. For tasks involving random radial masks at various compressed sensing (CS) ratios, meta-learning improved the average PSNR of T1 images by 1.54 dB and T2 images by 1.46 dB over the baseline. Even when tested on unseen CS ratios, meta-learning achieved further gains of 1.69 dB for T1 and 1.12 dB for T2 images. For random Cartesian masks with multiple undersampling rates, the T2 image PSNR improved by up to 1.87 dB compared to the conventional method.
The current study builds upon this foundation by extending the meta-learning paradigm to multi-modality and multi-coil reconstruction, with a provably convergent algorithm and a unified fusion-reconstruction[26] framework. The most relevant work[27] proposed an optimization-based meta-learning model that generalizes across sampling patterns and anatomical structures. Their framework integrates model-based and learning-based components and introduces meta-knowledge adaptation for unseen acquisition settings.
5.1 result
The test dataset from the fastMRI https://fastmri.med.nyu.edu/ provides 15 channel undersampled k-space data, the sampling is Kartesian Equispaced mask with 4-fold acceleration and 8-fold acceleration. I have checked, 4-fold acceleration is the same mask as what I used in the experiment. (VN provided their data set also released from NYU, but I’m not sure weather these data are the same as VN published.) They provide multi-coil coronal proton-density without fat suppression ground truth knee images. (Same type of image as we used.) The test data set is too large, I randomly choose 3 successive k-space multi-coil images (apply our mask) in 10 different sequences to test, so I test 30 data, here, "error" is relative error. The testing weights used from our experiment is: Phase5, Avg REC PSNR is 41.7114 dB, Avg relative error is 0.026917, Avg ssim is 0.9719. (This is the result of our testing data set.)
5.1.1 Comparing masks

### v2

# Task-Adaptive Reconstruction for Undersampled Multi-Coil MRI via Optimization-Guided Meta Learning

@@ Abstract
We propose a unified deep meta-learning framework for accelerated magnetic resonance imaging (MRI) that jointly addresses multi-coil reconstruction and cross-modality synthesis. Motivated by the limitations of conventional methods in handling undersampled data and missing modalities, we introduce a task-adaptive optimization network where meta-learned parameters control the reconstruction dynamics across varying acquisition conditions. Each phase of the network mimics a step of an adaptive forward-backward scheme with extrapolation, enabling the model to incorporate both data fidelity and nonconvex regularization in a principled manner. To enhance generalization across different acquisition settings, we integrate meta-learning, which enables the model to rapidly adapt to unseen sampling patterns and modality combinations using task-specific meta-knowledge. The proposed method is evaluated on the open source datasets, showing significant improvements in PSNR and SSIM over conventional supervised learning, especially under aggressive undersampling and domain shifts. Our results demonstrate the synergy of unrolled optimization, task-aware meta-learning, and modality fusion, offering a scalable and generalizable solution for real-world clinical MRI reconstruction.
@@ 1 Introduction
MRI reconstruction from undersampled measurements remains a challenging inverse problem. While parallel imaging and compressed sensing have significantly reduced acquisition time, aggressive undersampling still leads to aliasing artifacts and unstable reconstruction. Recent deep learning approaches have shown strong performance but often require retraining when acquisition conditions change. Despite these advantages, MRI examinations often require long acquisition times, which may lead to patient discomfort, motion artifacts, and limited scanner throughput. To mitigate these limitations, accelerated acquisition techniques have been developed to reduce the amount of measured data while maintaining acceptable reconstruction quality.
Parallel Imaging is widely used to accelerate imaging by employing multiple receiver coils to acquire signals simultaneously. Each coil observes the underlying image through a spatial sensitivity profile, allowing partial sampling of k-space while preserving sufficient information for reconstruction. However, recovering a high-quality image from undersampled measurements remains an ill-posed inverse problem. The reconstruction process must balance fidelity to the acquired data with prior knowledge about the structure of medical images.
In parallel with advances in reconstruction techniques, clinical MRI increasingly relies on multi-modality imaging. Different MRI contrasts, such as T1-weighted, T2-weighted, and FLAIR images, provide complementary information about tissue composition and pathology. Combining information from multiple modalities can significantly improve diagnostic interpretation. However, acquiring all modalities for every patient is often impractical due to time constraints, patient motion, and acquisition artifacts. Consequently, there is growing interest in methods that can jointly reconstruct undersampled data and infer missing modalities.
Deep learning has recently shown strong potential for solving inverse problems in medical imaging. In particular, algorithm-unrolled neural networks have emerged as a promising framework for integrating domain knowledge with data-driven learning. These models interpret iterative optimization algorithms as neural network architectures, where each iteration corresponds to a learnable computational block. This structure allows the model to maintain physical consistency with the imaging process while learning effective image priors from data.
Although unrolled networks have achieved encouraging results in MRI reconstruction, two limitations remain. First, many existing approaches focus solely on single-modality reconstruction and do not explicitly exploit relationships between different imaging contrasts. Second, models trained for a specific sampling pattern or acquisition protocol often struggle to generalize to new scanning conditions.
To address these challenges, this work introduces a unified reconstruction framework that combines algorithm unrolling with meta-learning. The proposed approach simultaneously handles multi-coil reconstruction and cross-modality synthesis from undersampled measurements. By incorporating meta-learning into the reconstruction process, the model learns transferable knowledge that enables rapid adaptation to new acquisition patterns and modality configurations. In addition, the underlying optimization algorithm provides theoretical convergence guarantees, ensuring stability of the iterative reconstruction process.
The main contributions of this work can be summarized as follows:
1. Unified Reconstruction and Modality Synthesis.
We propose a framework that jointly reconstructs multi-coil MR images and synthesizes missing modalities within a single optimization-driven architecture.
2. Algorithm-Unrolled Network with Adaptive Optimization.
The proposed model is derived from an adaptive forward–backward optimization scheme with extrapolation, enabling each network phase to correspond to a principled optimization step.
3. Meta-Learning for Cross-Task Adaptation.
A meta-learning mechanism is introduced to improve generalization across varying sampling patterns, acceleration factors, and modality combinations.
4. Convergence-Aware Network Design.
The reconstruction process is guided by a provably convergent optimization algorithm, providing theoretical support for the stability of the network.
Experimental results on publicly available datasets demonstrate that the proposed framework achieves improved reconstruction accuracy and robustness compared with conventional supervised learning approaches, particularly under aggressive undersampling and domain shifts.
@@ 2 Related Work
More recently, a diffusion model for accelerated MRI[8] and quantitative MRI [3]reconstruction using domain-conditioned prior guidance. This method leverages generative modeling to learn sample-consistent priors conditioned on domain information, achieving state-of-the-art results in both structural fidelity and quantitative accuracy.
@@ 3 Method
The goal of this work is to reconstruct high-quality MRI images and synthesize missing modalities from undersampled multi-coil k-space measurements. To achieve this objective, we design a learning framework derived from an optimization algorithm and implement it as a deep unrolled neural network.
Let $x$ denote the set of reconstructed images corresponding to different modalities, and let the observed measurements be obtained from multiple receiver coils under an undersampling pattern. Unlike conventional reconstruction models that employ a fixed regularization term, our formulation allows the regularization component to be dynamically adapted through meta-learned parameters. This enables the reconstruction process to adjust its implicit prior based on the acquisition task.
The detailed steps are described in Algorithm 1.
1:Input: Initial estimate $x^{(0)}$, smoothing parameter $\eta^{(0)}>0$, extrapolation factor $\tau^{(0)}\in(0,1)$, decay constants $\sigma,\mu,\delta\in(0,1)$, stopping threshold $\epsilon>0$
2:for $k=0,1,\dots,K-1$ do
3:  Compute a fidelity-driven update
$z^{(k)}=x^{(k)}-\alpha_{k}\nabla f\!\left(x^{(k)}\right).$
4:  Compute a regularization-aware candidate
$u^{(k)}=z^{(k)}-\beta_{k}\nabla r_{\eta^{(k)}}\!\left(z^{(k)}\right).$
5:  if $\phi_{\eta^{(k)}}(u^{(k)})\leq\phi_{\eta^{(k)}}(x^{(k)})-\frac{\delta}{2}\|u^{(k)}-x^{(k)}\|^{2}$ then
6:   Set $\bar{x}^{(k+1)}=u^{(k)}$
7:  else
8:   Set
$\bar{x}^{(k+1)}=x^{(k)}-\gamma_{k}\nabla\phi_{\eta^{(k)}}\!\left(x^{(k)}\right).$
9:  end if
10:  Form an extrapolated candidate
$y^{(k+1)}=\bar{x}^{(k+1)}+\tau^{(k)}\big(\bar{x}^{(k+1)}-x^{(k)}\big).$
11:  if $\phi_{\eta^{(k)}}(y^{(k+1)})\leq\phi_{\eta^{(k)}}(\bar{x}^{(k+1)})$ then
12:   Accept $x^{(k+1)}=y^{(k+1)}$ and update
$\tau^{(k+1)}=\min\!\left\{\frac{\tau^{(k)}}{\mu},\,1\right\}$
13:  else
14:   Accept $x^{(k+1)}=\bar{x}^{(k+1)}$ and update
$\tau^{(k+1)}=\mu\tau^{(k)}.$
15:  end if
16:  if $\|\nabla\phi_{\eta^{(k)}}(x^{(k+1)})\|<\sigma\eta^{(k)}$ then
17:   Set $\eta^{(k+1)}=\sigma\eta^{(k)}$
18:  else
19:   Set $\eta^{(k+1)}=\eta^{(k)}$
20:  end if
21:  if $\eta^{(k+1)}<\epsilon$ then
22:   break
23:  end if
24:end for
25:Output: $x^{(k+1)}$
Algorithm 1 Adaptive Forward–Backward Reconstruction Scheme
@@
The training objective is formulated as a composite loss that jointly penalizes synthesis error, perceptual inconsistency, cross-domain mapping discrepancy, and reconstruction error. The loss is defined as
$\mathcal{L}(x_{0},x_{1};\hat{x}_{0},\hat{x}_{1})=$ $\;\|x_{0}-\hat{x}_{0}\|_{2}^{2}$ (1a)
$+\lambda_{1}\big(1-\mathrm{SSIM}(x_{0},\hat{x}_{0})\big)$ (1b)
$+\lambda_{2}\left\|\tilde{\mathbf{g}}_{0}\!\left(\mathcal{G}\big(\mathbf{g}_{1}(x_{1})\big)\right)-\hat{x}_{0}\right\|_{2}^{2}$ (1c)
$+\lambda_{3}\|x_{1}-\hat{x}_{1}\|_{2}^{2}.$ (1d)
Here, $x_{0}$ denotes the synthesized image produced by the synthesis branch, while $x_{1}$ represents the reconstructed image obtained at the $T$-th phase of the network. The corresponding reference images are denoted by $\hat{x}_{0}$ and $\hat{x}_{1}$.
The loss incorporates four complementary terms. The first component measures the squared reconstruction error between the synthesized image and its reference. The second term introduces the structural similarity index (SSIM) to preserve perceptual structure and contrast information. The third component enforces a cycle-style consistency constraint by mapping the reconstructed image through the learned transformation $\tilde{\mathbf{g}}_{0}(\mathcal{G}(\mathbf{g}_{1}(\cdot)))$ and comparing the resulting estimate with the synthesized target. The final term penalizes the deviation between the reconstructed output and its ground-truth image. The scalar weights $\lambda_{1},\lambda_{2},\lambda_{3}$ control the relative influence of the perceptual, consistency, and reconstruction penalties during training.